OpenStudy (anonymous):

given g(x)=square root of x^2 =4 and f(x)=1/x^2 find (f x(times) g)(x)

OpenStudy (raffle_snaffle):

\[g(x)=\sqrt{x^2} and f(x)=1/x^2\]

OpenStudy (raffle_snaffle):